Use the future value formula to find the indicated value.


FV= 10,000; i= 0.03; PMT= $200; n= ?

n= ( Round up to the nearest integer as​ needed.)
Answer #1

Solution:-
Formula Future value of annuity (FV) = PMT * {(1 + i)n – 1}/i

FV = 10,000 ,i = 0.03 PMT = 200

=> 10000 = 200 * (1.03n – 1)/0.03

=> (1.03n – 1)/0.03 = 10000 / 200 = 50

=> (1.03n – 1) = 50 * 0.03 = 1.5

=> 1.03n = 2.5

=> log1.03n = log 2.5

=> nlog1.03 = log 2.5

=> n = log 2.5/log 1.03

=> n = 0.3979/0.0128

=> n = 31.085

n = 31